A user starts to download a 60-second movie encoded at 3Mbps. The Internet connection provides an average data rate of 1 Mbps.

A) How long will it be before the movie starts to play in the following cases?

1. the video application downloads the video before watching it. (the application plays the video once it is downloaded).

2. The video application uses progressive download.

B) How would your answer change if the average rate is 3Mbps?​
Answer #1

As movie is of 60sec encoded at 3Mbps so total size of movie is 60x3=180Mb

A. 1. As in this case firstly whole movie will download then it will play ...so time=180/1=180sec or 3min.

2.As in this progressive download is followed so video will start after 3 sec...because in 3seconds 1st second of video will load itself as video is encoded at 3Mbps and we are having speed of 1Mbps

B. 1. As in this case firstly whole movie will download then it will play ...so time=180/3=60sec or 1min.

2.As in this progressive download is followed so video will start after 1 sec...because in 1sec 1st second of video will load itself as video is encoded at 3Mbps and we are also having speed of 3Mbps